OpenStudy (anonymous):

Please check my solution: Find the absolute value of 3+4x My solution: -(3+4x) when x < 0 +(3+4x) when x >= 0

ganeshie8 (ganeshie8):

|3+4x|

ganeshie8 (ganeshie8):

find its vertex,

ganeshie8 (ganeshie8):

vertex = (-3/4, 0)

ganeshie8 (ganeshie8):

so, your solution should be :- -(3+4x) when x < -3/4 +(3+4x) when x >= -3/4

ganeshie8 (ganeshie8):

see if that makes some sense

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

Got it! I saw your answer, and tried to do it myself. 3+4x becomes: 3+4x or -3-4x 3+4x = 0 4x = -3 x = 3/4 -3-4x = 0 -4x = 3 x = -3/4

ganeshie8 (ganeshie8):

^ exactly ! -3/4 is the turning point

OpenStudy (anonymous):

Thanks! :)

ganeshie8 (ganeshie8):

np :)